# Knowledge Prompting: How Knowledge Engineers Use Large Language Models

### v1

@@ 5. Discussion
In this section we consider some of the key issues that arose from our results. The hackathon consisted of $77$ explorations of how LLMS can be used to support KE. Presenting the results of each of these explorations in detail is outside the scope of this paper, but there are some key overarching issues for their use that we present here. We then explore in some detail issues around skills, in particular prompting and bias detection and mitigation. Finally, we highlight the importance of transparency and explainability in KGs and we introduce the concept of “KG cards”.
5.1. Issues in using large language models for knowledge engineering
A fundamental conflict occurs in the use of LLMs in KG work because of the ambiguity in the output of generative AIs. This ambiguity lies mostly in the inherent irreproducibility of results garnered using generative AI. This means that contributions made by LLMs, such as the creation of relations or the similarity between terms, are based on statistical probability rather than consistent fact. This lack of consistency is also problematic for scientific methods. This is consistent with previous studies arguing that knowledge extraction remains challenging despite the growing field of NLP (Noy et al., 2019; Allen et al., 2023). However, one notable use of the LLMs for KE is to potentially create a structured dataset from a variety of unstructured data i.e. for data integration purposes. Other research (Walker et al., 2023) has found that given a prompt that includes the term “dataset”, in the absence of a dataset, ChatGPT may attempt to “curate” a dataset like an approximation from unstructured data. Therefore, this application seems highly plausible for KE and, particularly, knowledge extraction in a structured format which is familiar to existing practices. Following the usual actions of cleaning and transforming, the dataset can be used for KG construction. Yet, the accuracy of LLM results remains an issue.
@@ 5.1. Issues in using large language models for knowledge engineering
Following the knowledge extraction difficulties, another particular area of concern was evaluation. Attempting to evaluate LLM support fro KE tasks would simply exacerbate existing evaluation challenges, which were many. These include evaluating accuracy (i.e., syntactic and semantic) and coherency (i.e. consistency and validity), emphasising that we may assess accuracy automatically, but most of the time, coherency requires manual interventions. Furthermore, concerns about coverage (i.e. completeness and representativeness ) and succinctness (i.e. conciseness and understandability) are also relevant to input as well as output, and should be included as a consideration. This is a similar problem to that experienced in crowd-sourced KGs such as Wikidata (Abián et al., 2017), where variations in the input (such as background and skill of contributors) also demands evaluation as much as the output.
Regarding the evaluation of LLM outputs, one interesting suggestion was made related to using adversarial algorithms. “Would something or a product like that stand the scrutiny of an adversarial network, who is trying to maybe test or break the new information derived by the existing system. So can there be a counter point counter punch that tries to break it, and if it cannot be broken or falsified?” Recently, an adversarial attack on BARD (Google, 2024) asking to repeat one iteratively one word eventually got the LLM to reveal gigabites of its training data, demonstrating the usefulness of such processes in testing systems (Nasr et al., 2023).
@@
Another proposed aspect of evaluation was human satisfaction. Other research has shown that the natural language element of LLMs, often read by humans as possessed of emotion, can make people feel that the process of using the LLM was successful even if the end result was not (Walker et al., 2023). This could be potentially problematic for KE tasks, which requires logic over feelings. A comparable, well-researched situation is that of emotion in financial markets, which can move prices up or down based on public mood rather than underlying economic fundamentals
5.2. Key skills - prompting
@@
However, ultimately, prompting was utilised for all the different KE tasks. This flags a new era and requirement for a new skill set in the field of KE. This suggests that the creation of templates (using a method such as that suggested in (Liu et al., 2023b)) for the KE tasks could benefit the field towards KG construction and fill in gaps in prompting skill level. This has parallels in the introduction of crowd-sourcing to KG construction around a decade ago. Using crowdsourcing means for large-scale KG construction was a successful innovation supporting many intelligent applications today. Additionally, incorporating a prompt engineer in the process, the “knowledge prompting engineer” could also bridge the skill gaps and enhance performance.
5.3. Key skills - bias
@@
There was a distinct lack of engagement with the processes of bias mitigation. This tended to arise from limited exposure to techniques or theories for identifying bias in the model. While sometimes this was due to a lack of opportunity for KEs to do this, in other cases, it was because of a lack of interest. It is difficult to argue that the identification and eradication of bias should be left to a specific set of experts, rather than integrating core safety facets into the training of those responsible for building AI. Given the variety of safety and ethical aspects and potential harms in AI, it may be that space for “ethical deliberation” in the KG construction process is more valuable than simply providing, for instance, a list of biases for engineers to look out for (Gogoll et al., 2021).
Our results illustrated a data-centric approach to bias in KGs, suggesting that the source of bias was more likely to be the data than the graph model itself. However, a few recognised that KG/Ontology/LLM is also important and the data may well be secondary. Including LLMs in the construction of KGs may complicate these sources of bias further and require enhanced awareness (Huang et al., 2023).
@@
The integration of LLMs into the KE processes may additionally create further potential for harm, such as when, intentionally or unintentionally, the LLM can be manipulated to overcome privacy (as shown in the BARD example above (Nasr et al., 2023)). This suggests an integrated review of the possible sources of harms introduced to KG when LLMs are utilised would be valuable.
5.4. KG cards to reduce harms
One way to implement this review of harms would be to develop KG data cards (Mitchell et al., 2019; Pushkarna et al., 2022), a set of detailed descriptions for transparency and explainability to ensure we produce and use trustworthy KGs. Considering the data cards model (Pushkarna et al., 2022), a set of similar content themes could be used to guide researchers working on KG construction to supplement their datasets, in this case, KGs, with explanations related to provenance, coverage etc. Given the lack of training in ethics and, therefore, the need for greater guidance in this area, KG cards can support extensive documentation and guide researchers in order to avoid common safety issues and mistakes.
Data cards include a list of $3131$ content themes related to publishers, funding, licence, versions, features, sources, motivation, applications, limitations, cleaning, labelling, descriptive statistics, fairness, and term/domain definitions (Pushkarna et al., 2022). Inspired by these themes in Figure 1, we suggest a list of “KG cards” with main themes provenance, construction, knowledge, completeness, safety and licencing. With provenance, we suggest tests similar to Holistic Evaluation of Language Models (HELM) (Bommasani et al., 2023) for language model transparency and data cards (Pushkarna et al., 2022) descriptions for dataset explainability. For construction, we propose details related to the structure of the graph for transparency in data processing, and for knowledge, we seek to describe the domain, the different languages, and the range of knowledge included (e.g. the list of UK prime ministers up to $20232023$).
Heist et al. (Heist et al., 2020) presented an overview and comparison of publicly available KGs. Using his recommendations, for the completeness card, we suggest providing details similar to network analysis. These are information about the number of nodes, relations, and average linkage, as well as about class hierarchy. This can offer and understanding for the KG coverage and potentially can support the next card about safety tests.
@@ 5.4. KG cards to reduce harms
While data cards describe provenance in detail, there is a lack of focus on aspects of safety, particularly bias. “Fairness” as a concept is vague, culturally dependent and difficult to operationalise (Munn, 2023; Dodds, 2020). For KG cards, we suggest a set of safety checks related to data, the application of the KG, KG relations, fairness and explainability. With this we aim to provide explanations and support for bias and ethics.
Finally, licensing for the data used to build the graph and for the result graph can support its reuse.
Figure 1. Six suggested KG cards with related details.
@@
Using well-structured cards for transparency and explainability motivated others with the Hugging Face platform, inspired by model cards (Mitchell et al., 2019), to create a template called dataset cards for its shared datasets. [^11] https://huggingface.co/docs/hub/datasets-cards The attempts from essential platforms in the field of AI, like Hugging Face, to create frameworks for explainable datasets show the importance of data provenance and process in the AI era. Table 4 shows an alignment between Hugging Face’s dataset cards and our suggested KG cards. Most cards match different needs and directions depending on the product (i.r. dataset or KG); however, we suggest incorporating state-of-the-art metrics and emphasising the importance of ethical issues.
@@
Table 4. Alignment between KG cards, and the Hugging Face dataset cards.
KG Cards Hugging Face Dataset Cards
Provenance Dataset Creation
Construction Dataset Creation
Knowledge Dataset Description
Completeness Dataset Structure
Safety Considerations of Using the Data
Licensing Additional Information
@@
In addition, Hugging Face has also adapted the set of model cards for transparency and detailed documentation of their models. [^12] https://huggingface.co/docs/hub/model-cards We suggest a similar tactic for the explainability of the KG embedding model in the KE field. The set of model cards (Mitchell et al., 2019) can be used to support documentation with a metadata template and enhance model transparency. Table 5 presents the list of model cards with descriptions and one example of the KG embedding model TransE. TransE, which stands for Translating Embeddings for Modeling Multi-relational Data, is a state-of-the-art method for creating KG embeddings. Building upon the origin TransE model, more popular models have been created, like TransR (rotating) and TransF (folding) (Dai et al., 2020). The TransE example in the table shows that KG embedding model studies offer detailed documentation for model training and evaluation, but the lack of demographic details and ethical considerations flag the need, similar to datasets, to adapt specialised instructions for model documentation in the field of KE.
@@
Table 5. This Table presents the model cards by Michelle et al. (Mitchell et al., 2019), with descriptions and the alignment for one example of the KG embeddings model TransE.
Model Cards Description TransE
@@
Model details Basic information about the model Bordes et al. (Bordes et al., 2013) 2013 Translation-based model {bordesan, nusunier, agarciad}@utc.fr {jweston, oksana}@google.com
@@
Intended use Use cases that were envisioned during development Link prediction (details in (Bordes et al., 2013) section 4.3)
@@
Metrics Metrics should be chosen to reflect potential real-world impacts of the model Head and Tail example (details in (Bordes et al., 2013) Table 5)
Evaluation data Details on the dataset(s) used for the quantitative analyses in the card WordNet (Fellbaum, 2010) and Freebase (Bollacker et al., 2008) (details in (Bordes et al., 2013) section 4.1)
Training data May not be possible to provide in practice. When possible, this section should mirror Evaluation Data. If such detail is not possible, minimal allowable information should be provided here, such as details of the distribution over various factors in the training datasets Trained on a large-scale split of Freebase containing 1M entities, 25k relationships and more than 17M training samples (details in (Bordes et al., 2013) section 4.1)
Quantitative analysis Unitary and intersectional results Mean and Hits metrics, compered with RESCAL (Nickel et al., 2011), SE (Bordes et al., 2011), SME(Bordes et al., 2014), LFM (Jenatton et al., 2012) (details in (Bordes et al., 2013) section 4.2)
@@
6. Conclusion
LLMs herald the emergence of a new era for many fields. It is inevitable that they will change KE in many ways. In what direction this will lead us is not clear. The hype around conversational generative AI is such that may one day render KGs obsolete. Currently the volatility and unreliability of generative AI means that KGs will continue to be a trusted source, although it is impossible to tell if this will still be true if LLMs become more reliable in the future. Today, however, the added efficiency in terms of addressing volume based tasks is of great potential for use alongside more traditional methods. In terms of adoption of new ways of creating KGs this recent development is somewhat analogous to the introduction of crowd sourcing to KE; an innovation that has subsequently proved of immense value. However, it is crucial that those working in the field are appropriately skilled and supported to work with LLMs in an effective, productive and safe fashion. As well as training, this implies the need for documentation to accompany LLM-assisted KGs such as prompt templates and KG cards.
6.1. Limitations
Participants were selected from European research labs that have a strong profile in either publishing in KE venues, or in offering well-used KE industry products. Given the number of participants not all KE tasks were investigated, the experiments were relatively small scale and groups worked for only three days on the tasks. Furthermore, most interviewees were different-level PhD students, missing the opinion of researchers with more than $1010$ years of experience in the field.
6.2. Contributions and Future Work
This is one of the first attempts to understand the interaction between knowledge engineers and LLMs. We have identified and presented key areas of strength and challenge for KE working with LLMs, and we have built on the suggestion of implementing data cards for KGs and propose extending these to more comprehensively address safety within KGs developed with LLMs.
Future work may look at how LLMs can support less expert users in developing KGs, as this has been a key use of lLMs in other fields such as coding, report writing, etc. In particular, a more detailed study on practical methods for identifying and mitigating against bias in LLM-assisted KGs is required, across the entire development pipeline.
6.3. Acknowledgements

### v3

@@ 5 Discussion
In this section, we consider some of the key issues that arose from our results. We examine in more depth issues related to skills, in particular prompting and bias detection and mitigation, and conclude by emphasising the importance of transparency and explainability in KGs. Finally, we introduce the concept of “KG Cards” and we suggest the use of the existing Model Cards for the KG embedding models.
5.1 Issues in Using Generative AI for Knowledge Engineering
A fundamental challenge in using generative AI for KE stems from the ambiguity and variability of its outputs. This is most evident in the inherent irreproducibility of results from single, directly prompted LLMs, where generated relations or term similarities are based on statistical probability rather than consistent fact. Such inconsistency undermines scientific rigour and complicates error tracing, making control and refinement difficult. (Since we conducted the interviews, this has been at least partly mitigated with retrieval agumented generation (RAG) [90], which enables LLMs to retrieve relevant documents and integrate their information to user prompts before providing query responses; however, issues of bias and fairness around the integration of specific external knowledge sources and the models themselves persist [80, 179].) These concerns align with prior findings that knowledge extraction remains challenging despite advances in NLP [8, 119]. Compared to pre-LLM human-in-the-loop KE tools [49], LLM-assisted approaches offer more end-to-end workflows and streamlined human–machine interaction, but also obscure the origin of errors. However, recent advances—such as tool calling [141], retrieval-augmented generation [90, 50], confidence elicitation [180, 52], and self-assessment [142]—have significantly mitigated these issues. These developments could lead to more reliable, and reproducible LLM-assisted KE workflows, thus change the attitude of knowledge engineers towards LLM-assisted systems.
Another notable application of LLMs in knowledge engineering is their potential to generate structured datasets from various unstructured data sources, particularly for data integration purposes. Research has shown that, when given a prompt containing the term “dataset”, in the absence of a dataset, ChatGPT may attempt to “curate” a dataset like an approximation from unstructured data [173]. Therefore, this application seems highly plausible for KE and, particularly, knowledge extraction in a structured format which is familiar to existing practices. Following the usual actions of cleaning and transforming, the dataset can be used for KG construction.Recent advances in the structured generation capabilities of LLMs have further strengthened their potential for such applications [83, 178, 39, 98]. Yet, the accuracy of LLM results remains an issue.
@@ 5.1 Issues in Using Generative AI for Knowledge Engineering
Following the knowledge extraction and integration, another particular area of concern was evaluation. Attempting to evaluate LLM support for KE tasks would simply exacerbate existing evaluation challenges, which were many. These include evaluating accuracy (i.e., syntactic and semantic) and coherency (i.e. consistency and validity), emphasising that we may assess accuracy automatically, but most of the time, coherency requires manual interventions. Furthermore, concerns about coverage (i.e. completeness and representativeness ) and succinctness (i.e. conciseness and understandability) are also relevant to input as well as output, and should be included as a consideration. This is a similar problem to that experienced in crowd-sourced KGs such as Wikidata [1], where variations in the input (such as background and skill of contributors) also demands evaluation as much as the output. Recently, the community has devoted increased attention to evaluation, emphasising both the development of robust evaluation benchmarks and the adoption of human-centered principles, strategies, and paradigms [190, 101, 97, 65, 63]. Recent research has explored the use of LLMs as a judge to evaluate the results of competency questions, but suggests these are still at the stage of being useful as supporting human inputs, rather than as trustworthy standalone tools [82]. Regarding the evaluation of generative AI outputs, one interesting suggestion was made related to using adversarial algorithms. “Would something or a product like that stand the scrutiny of an adversarial network, who is trying to maybe test or break the new information derived by the existing system. So can there be a counter point counter punch that tries to break it, and if it cannot be broken or falsified?” An adversarial attack on BARD [57] asking to repeat one iteratively one word eventually got the model to reveal gigabits of its training data, demonstrating the usefulness of such processes in testing systems [115].
@@
Another proposed aspect of evaluation was human satisfaction. Other research has shown that the natural language element of LLMs, often read by humans as possessed of emotion, can make people feel that the process of using the model was successful even if the end result was not [173]. This could be potentially problematic for KE tasks, which requires logic over feelings. A comparable, well-researched situation is that of emotion in financial markets, which can move prices up or down based on public mood rather than underlying economic fundamentals.
5.2 Key Skills for Effective Prompting
@@
However, ultimately, prompting was utilised for all the different KE tasks. This flags a new era and requirement for a new skill set in the field of KE. This suggests that the creation of templates (e.g., [6, 99, 193]) for the KE tasks could benefit the field towards KG construction and fill in gaps in prompting skill level. This has parallels in the introduction of crowd-sourcing to KG construction around a decade ago. Using crowdsourcing means for large-scale KG construction was a successful innovation supporting many intelligent applications today [2, 143, 84]. Furthermore, leveraging (semi-)automatic prompting methods [183] and incorporating domain-specific prompting strategies [67] can help address skill gaps and improve performance. Very recent research suggests that although reasoning has received a lot of attention as a prompting strategy, for some major LLMs this does not provide substantial improvement over simple instruction and task demonstration for knowledge extraction from text [135]. This implies that in this area at least, proficiency in prompting may easily be acquired.
5.3 Key Skills to Mitigate Bias
@@
Knowledge engineers and domain experts involved in generative AI-supported KE projects are key stakeholders in promoting responsible AI practices. Ensuring that AI systems are developed, assessed, and deployed in a safe, trustworthy, and ethical manner is critical for fostering appropriate trust among these professionals. There was a distinct lack of engagement with the processes of bias mitigation. This tended to arise from limited exposure to techniques or theories for identifying bias in the model. While this was sometimes due to a lack of opportunity for knowledge engineers to acquire capabilities, in other cases it was due to a lack of a sense of responsibility for bias mitigation. It is difficult to argue that the identification and eradication of bias should be left to a specific set of experts, rather than integrating core safety facets into the training of those responsible for building AI. Given the variety of safety and ethical aspects and potential harms in AI, it may be that space for “ethical deliberation” in the KE process is more valuable than simply providing, for instance, a list of biases for engineers to look out for [56].
While capacity building (training) and bringing awareness to the importance of taking responsibility for ownership of AI safety issues appear to be two separate issues, they are contingent upon each other and should be treated as two sides of the same coin. Proper exposure to bias awareness and mitigation techniques should include training on when and where to review aspects of safety – thus integrating the idea of responsibility for safety with knowledge of the actions to take.
a
@@
Our results illustrated a data-centric approach to bias in KE, where our participants felt that the source of bias was more likely to be the data than the graph model itself, and relatively few recognised that the model may be a source of bias. Including generative AI in the construction of KGs will introduce a further source of bias [69]. The integration of generative AI into the KE processes may additionally create further potential for harm, such as when, intentionally or unintentionally, the model can be manipulated to overcome privacy (as shown in the adversarial attack example above [115]). This suggests an integrated review of the possible sources of harms introduced to KE when generative AI models are utilised would be valuable.
A number of the issues raised by participants in the hackathon - particularly those around evaluation and ethics – speak to deeper, more sustained challenges within the discipline. Our findings show that many knowledge engineers feel ethical issues are the responsibility of ’someone else, and that bias mainly resides in data and not in any part of model construction. This finding is similar to other studies with AI engineers or practitioners, who feel either that the responsibility for mitigating harm does not lie with them, or that as individuals they cannot do much about it [87]. A great deal of recent work on LLMs for KE emphasises that research into bias in this area is of great importance [48, 130, 9, 54, 191]. However, no study has directly addressed bias and fairness in LLM-based KG construction and there are no concrete proposals on how to address it. This suggests that work needs to be done to embed a greater understanding of where bias lies, and how KE can address it. LLMs are not neutral and have bias up and down the chain that must be understood and appraised [80, 179]. Benchmarks for evaluating LLM-generated ontologies have been proposed [134], but are limited to assessing the presence/absence of classes and properties and their level of abstraction, not their possible issues around bias and fairness. A starting point would be measuring bias and fairness in LLM-based KG outputs, for which inspiration could be taken from concepts such as Vendi (diversity evaluation) scores and the multii-metric HELM approach [47, 93].
5.4 KG Cards to Enhance Transparency
As described in Section 2.3, prior research aimed at improving transparency and explainability has introduced frameworks such as Model Cards [109], Data Cards [138], and Data Statements [15]. Implementing our review of harms we point out the need for similar practices for KGs. The aforementioned Cards can not fully serve the KG’s needs because they do not consider the unique design aspects of KGs. These Cards fail to describe the ontology and the KG data, such as relations and schemas. Building on these studies, we propose the concept of Knowledge Graph Cards (KG Cards) as a documentation framework to promote transparency and explainability, ensuring the creation and use of trustworthy KGs. Drawing inspiration from the flexibility and extensibility of Data Cards, and considering the foundational representation format of most KGs, KG Cards are designed to enhance transparency and trustworthiness by providing comprehensive and structured documentation.
Figure 2: Six suggested sections of KG Cards with related details.
As shown in Figure 2. It contains 6 main themes:
  • Knowledge. The Knowledge section aims to provide an overview of the domain, the languages represented, and the range of knowledge included. It may also document ontology development artifacts, such as competency questions, which help define the knowledge graph’s (KG) scope and support ontology testing.
  • Provenance. The Provenance section is intended to capture detailed provenance information about the KG. Various platforms and frameworks, such as the W3C Provenance Ontology (PROV-O) [88], have been proposed for managing provenance in KGs [81]. Additional provenance details can be documented here, including information about crowdsourcing (e.g., platform used, demographics of crowdworkers, and domain experts), references to datasets (and Data Cards, where applicable) used in KG creation, and other contextual information.
  • Graph Construction. The Graph Construction section outlines the details of the data preprocessing methods and the approaches used during the construction and curation process. This includes techniques for knowledge extraction, integration, and the use of KG embedding models.
  • Quality. A crucial aspect of KG transparency and explainability comprises the quality assessment. Considering the KG quality assessment dimensions described in Section 2.1 and quality dimensions for linked data [188], we introduce the Quality section. This includes descriptions about accuracy, coverage, coherency, and succinctness. Many studies have focused on completeness [66, 127, 151, 70], which is also a vital dimension for explainability. Syntactic and semantic accuracy describes how errors are handled, while timeliness states the pace of updates. Coverage looks at the network details, stating statistics which can offer an understanding of the KG coverage and potentially can support the next card about safety tests. Moreover, coherency expresses contradictions in the data and succinctness describes the human-readable aspects, the URIs details, SPARQL query services, and community contact options.
  • Licensing. The Licensing section provides details about the licenses, publishers, funding sources, and associated policies.
@@ 5.4 KG Cards to Enhance Transparency
  • Transparency. While previous documentation methods provide detailed transparency principles, they often lack emphasis on fairness aspects. “Fairness” as a concept is vague, culturally dependent and difficult to operationalise [111, 38]. For KG Cards, we suggest the Transparency Card with a set of transparency and fairness checks related to data, the intended and unintended application of the KG, and explainability. With this, we aim to provide explanations and support for bias and ethics.
An example of the use of KG cards is presented in Appendix D, Table 8. The presented KG Cards represent a preliminary framework for enhancing transparency in KGs. Like previous model and dataset documentation methods, the proposed KG Cards are adaptable, allowing more sections and properties to be included based on the specific needs of the stakeholders in various contexts. Further research could offer valuable insights through extended analysis, helping to validate usability and identify potential modifications or enhancements to the structure.
@@
The use of well-structured documentation for transparency and explainability has inspired initiatives like the Hugging Face platform’s adoption of Dataset Cards and the Croissant metadata format for its shared datasets [5]. [^12] https://huggingface.co/docs/hub/datasets-cards Table 4 shows the alignment between Hugging Face’s Dataset Cards, Croissant metadata format, and our suggested KG Cards. Most cards match different needs and directions depending on the product (i.e., dataset or KG); however, we suggest incorporating state-of-the-art metrics and emphasising the importance of ethical issues.
@@
Table 4: Alignment between KG Cards, Hugging Face dataset cards, and Croissant metadata format.
KG Cards Hugging Face Dataset Cards Croissant Metadata Format
Knowledge Dataset Description The Dataset Metadata Layer
Provenance Dataset Creation The Resources Layer
Construction Dataset Creation The Structure Layer
Quality Dataset Structure The Dataset Metadata Layer
Licensing Additional Information The Dataset Metadata Layer
Transparency Considerations of Using the Data The Dataset Metadata Layer (RAI properties), The Semantic Layer
5.5 Model Cards to Reduce Harms
@@
Besides the Data Cards framework, Hugging Face has also adapted the set of Model Cards for transparency and detailed documentation of their models. [^13] https://huggingface.co/docs/hub/model-cards We suggest a similar tactic for the explainability of KG embedding models in the KE field. The set of Model Cards [109] can be used to support documentation with a metadata template and enhance model transparency. Table 5 presents the list of Model Cards with descriptions and one example of the KG embedding model, TransE [20]. TransE is a KG embedding model that models relationships in the graph by interpreting them as translations applied to the low-dimensional embeddings of entities. Inspired by the TransE model, several variant models have been developed, such as TransR (rotating) [95] and TransF (folding) [43]. The TransE example in the table shows that KG embedding model studies offer detailed documentation for model training and evaluation, but the lack of demographic details and ethical considerations flag the need, similar to datasets, to adapt specialised instructions for model documentation in the field of KE.
@@
Table 5: Model Cards by Michelle et al. [109], with descriptions and the alignment for one example of TransE model.
Model Cards Description Example Model (TransE)
@@
Model details Basic information about the model Bordes et al. [21] 2013 Translation-based model {bordesan, nusunier, agarciad}@utc.fr {jweston, oksana}@google.com
@@
Intended use Use cases that were envisioned during development Link prediction (details in [21] section 4.3)
@@
Metrics Metrics should be chosen to reflect potential real-world impacts of the model Head and Tail example (details in [21] Table 5)
Evaluation data Details on the dataset(s) used for the quantitative analyses in the card WordNet [42] and Freebase [18] (details in [21] section 4.1)
Training data May not be possible to provide in practice. When possible, this section should mirror Evaluation Data. If such detail is not possible, minimal allowable information should be provided here, such as details of the distribution over various factors in the training datasets Trained on a large-scale split of Freebase containing 1M entities, 25k relationships and more than 17M training samples (details in [21] section 4.1)
Quantitative analysis Unitary and intersectional results Mean and Hits metrics, compared with RESCAL [118], SE [22], SME[19], LFM [72] (details in [21] section 4.2)
@@
6 Conclusion
Generative AI heralds the emergence of a new era for many fields. It is inevitable that they will change KE in many ways. In what direction this will lead us is not clear. The hype around conversational generative AI is such that may one day render KGs obsolete. Currently the volatility and unreliability of generative AI means that KGs will continue to be a trusted source, although it is impossible to tell if this will still be true if LLMs become more reliable in the future. Today, however, the added efficiency in terms of addressing volume based tasks is of great potential for use alongside more traditional methods. In terms of adoption of new ways of creating KGs this recent development is somewhat analogous to the introduction of crowd sourcing to KE; an innovation that has subsequently proved of immense value. However, it is crucial that those working in the field are appropriately skilled and supported to work with generative AI in an effective, productive and safe fashion. As well as training, this implies the need for documentation to accompany LLM-assisted KGs such as prompt templates and KG cards.
6.1 Limitations
Participants were selected from European research labs that have a strong profile in either publishing in KE venues, or in offering well-used KE industry products. Given the number of participants not all KE tasks were investigated, the experiments were relatively small scale and groups worked for only three days on the tasks. Furthermore, most interviewees were different-level PhD students, missing the opinion of researchers with more than 10 years of experience in the field. It is worth mentioning as a limitation that there is a gap between the completion of data analysis and the publication of this study, which may influence the contemporaneity of the findings. However, the core insights remain robust and continue to provide meaningful contributions to how knowledge engineers use generative AI.
6.2 Contributions and Future Work
This is one of the first attempts to understand the interaction between knowledge engineers and generative AI. We have identified and presented key areas of strength and challenge for KE working with LLMs, and we have built on the suggestion of implementing Data Cards for KGs and propose extending these to more comprehensively address safety within KGs developed with generative AI.
Future work could explore how generative AI can solve other KE tasks and assistless experienced users in KE tasks, building on its established applications in other domains such as programming, report generation, and content creation. Additionally, further efforts are needed to deepen our understanding of the role generative AI plays in specific KE tasks and throughout the KG lifecycle. A particular focus could be placed on conducting comprehensive studies to identify practical strategies for detecting and mitigating biases in LLM-assisted KGs across the entire development pipeline. Key to this is understanding the scale of whether KE’s lack knowledge about safety, or whether the real barrier is obstacles to the implementation of this knowledge.
6.3 Acknowledgments